THE C_{2}N_{2} C^{1}B_{u} - X^{1}Σ_{g}^{+} TRANSITION: ASSIGNMENT OF THE EXCITED ELECTRONIC STATE, ITS ORIGIN AND VIBRATIONAL FREQUENCIES

\underline{JOSHUA B. HALPERN}, AND YUHUI HUANG, DEPARTMENT OF CHEMISTRY, HOWARD UNIVERSITY, WASHINGTON, DC 20059.

A new analysis is presented for the VUV C_{2}N_{2} C^{1}B_{u} - X^{1}Σ_{g}^{+} transition. Contrary to previous studies this is shown to be a linear to bent transition with a trans-bent C^{1}B_{u} excited electronic state. The electronic origin is at 59,848 cm^{−1}, about 550 cm^{−1}lower than previously thought ^{a}. Vibrational bands between 170 and 145 nm are assigned. The C^{1}B_{u} symmetric CN stretching frequency, ν_{1}, is as previously noted 2050 ¯+ 20 cm^{−1}. It has been possible to assign four of the five other vibrational frequencies. The symmetric CC stretching frequency, ν_{2}, is 926 ¯+ 20 cm^{−1}. The trans-bending vibrational frequency, ν_{3}, is 532 ¯+ 10 cm^{−1}, the torsional bending frequency, ν_{4}, is 110 ¯+ 10 cm^{−1}and the cis bending frequency, ν_{6}, is 140 ¯+ 10 cm^{−1}. Anharmonicities have also been determined for the various vibrational modes.

aS. Bell, et al., Mol. Spectrosc. 30 (1969) 162